How would I answer this problem, The width of the triangle is one-third of the length. If the perimeter is 40 cm, find the dimensions of the rectangle.
Found 2 solutions by Jk22, texttutoring:
Answer by Jk22(389)

the perimeter is two times width plus length, equals here two times length plus one third of length, which is two times one plus one third times length.
Hence we get forty equals eight third of the length, so
length equals three eighth of forty, which equals fifteen centimeter

Answer by texttutoring(324)

The perimeter formula is:

P = 2w + 2L

The width is: w = L/3

So take the w =L/3 and substitute it into the perimeter formula:

P = 2w + 2L
40 = 2(L/3) + 2L
40 = 2L/3 + 2L
40 = 2L/3 + 6L/3
40 = 8L/3
120 = 8L
L = 15

So length = 15 cm.
Width = Length/3
Width = 15/3 = 5 cm.